A mixture containing 6% boric acid is to be mixed with 4 liters of a mixture which is 25% boric acid in order to obtain a solution which is 12% boric acid. How much of the 6% solution must be used?
Answer by stanbon(75887)

A mixture containing 6% boric acid is to be mixed with 4 liters of a mixture which is 25% boric acid in order to obtain a solution which is 12% boric acid. How much of the 6% solution must be used?
---------------------------------
Acid EQUATIOn: 0.06x + 0.25*4 = 0.12(x+4)
6x + 100 = 12x + 48
6x = 52
x = 8 2/3 liters (amt. of 6% solution that must be added)
===============
Cheers,
Stan H.